OpenStudy (anonymous):

Graph the function y=x^2 + 2x - 3 Draw a chart to show at least three ordered pairs that are on the graph.

OpenStudy (anonymous):

Graph and some order pairs

OpenStudy (anonymous):

y = x2 - 2x - 3 is a parabola. Since the x2 is positive, it opens upward (concave-up). If you factor the right hand side, you get (x+1) (x-3) so that means that the x-intercepts are at -1 and +3. The vertex is halfway between these of course (at x = 1). That means the vertex is at x = 1 and y = 12 - 2(1) - 3 = -4. So draw a parabola shape with the vertex at (1, -4) and x intercepts at -1 and +3.

OpenStudy (anonymous):

Any questions?

OpenStudy (anonymous):

I understand it a bit, but i still have no clue what the ordered pairs would be.

OpenStudy (anonymous):

it should be (-4,5),(-3,0),(-2,-3),(-1,-4),(0,-3),(1,0),(2,5)

OpenStudy (anonymous):

^^^ lol

OpenStudy (anonymous):

there can only be 3 ordered pairs

OpenStudy (anonymous):

you can choose any three and just connect the dots

OpenStudy (anonymous):

(-1.4),(1,0),(2,5)

OpenStudy (anonymous):

all those order pairs are in the graph

OpenStudy (anonymous):

@chuy99 is right (y)

OpenStudy (anonymous):

alright I think I got it, thanks.

OpenStudy (anonymous):

yw @isaacG